Update, 20/10/2014, 09:29
Systems continue to function as normal, so we have marked Fridays issue as resolved.
For further information and updates from Friday please click below…
Update, 17/10/2014, 16:08
There continue to be a few teething problems with OWA access that are being worked on, but SLP continues to function as normal. This will be the last update for today from us, but there remains engineers working to finalise service restoration.
Update, 17/10/2014, 15:40
Engineers think that they have identified and fixed the issue… SLP currently appears to be functioning normally, but it will be monitored over the weekend.
For anyone interested the problem issue was to do with SSL on our front end TMG servers.
Update, 17/10/2014, 15:05
SWOne continue to work on resolving the issues with slow/intermittent access to SLP and OWA. Work is being undertaken on both servers and it is estimated that this may take up to 3 hours until which time when SLP/OWA will be unavailable.
If the restoration of the servers is unsuccessful then a manual rebuild and support from Microsoft will take place.
Every effort is being made to restore these services as soon as possible.
Colleagues need to be aware that at the present time it is difficult to predict or guarantee service level or availability during the coming weekend.
Update, 17/10/2014, 13:56
Two load balancing servers control the traffic to SLP, one of which is currently overloaded and not working at all well.
The second server is currently not working at all, and is being worked on by SWO engineers.
Since around 8:45am today (Friday 17th) the entire SLP appears to be unresponsive… SouthWest One are urgently looking at the problem and it should be back up again shortly.
9:24 – Various servers are currently being rebooted in an attempt to solve the problems.
9:24 – Many schools appear to also have email problems which may be related, engineers are investigating. UPDATE: Email issues have been resolved for those using Outlook. For those using Outlook Web Access there are still severe access problems.
9:49 – There are currently no further updates from SWO, there is still little or no service externally to SCC.
13:55 – One server is still overloaded and not working well, and one server is currently being recovered from backup in an attempt to get it working again. This may resolve the situation when complete. A Major Incident has been declared on this outage by SCC/SWO.
Apologies for this break in service which is outside of eLIM control… this notice will be updated as and when we have more information.